OpenStudy (anonymous):

Find the domain of the following functions: f(x)=(6x^2-2x+1)/3

OpenStudy (anonymous):

(- oo, oo)

OpenStudy (anonymous):

Domain is looking for numbers where the function is discontinuous.

OpenStudy (anonymous):

it is continuoues everywhere so its from negative infinity to infinity

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Yeah. I wasn't saying you were wrong, just trying to put out the general way of solving the problem without just giving the answer.

OpenStudy (anonymous):

oo yeah gotcha, i was giving a reason to my answer

OpenStudy (anonymous):

Alright, no worries.

OpenStudy (anonymous):

you guys are cool! very helpful i think i have a harder one

OpenStudy (anonymous):

And that is?

OpenStudy (anonymous):

f(x)=x+squartroot x^2+1

OpenStudy (anonymous):

cant figure out how to do the squartroot symbol

OpenStudy (anonymous):

Still looking for areas where the function is discontinuous. This one is a little different. Normally you check for values that make the square root bad. So that would be negative numbers in a square root.

OpenStudy (anonymous):

Here you would check \[\sqrt{x^2+1}\] and well what do you find out?

OpenStudy (anonymous):

since the x value will always be positive because it is squared it will be continuous everywhere aswell

OpenStudy (anonymous):

hmm what if its sqrt root of x^2-1 instead. will it be same

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

x cant be 1

OpenStudy (anonymous):

Not quite. If x=1 it's still valid.

OpenStudy (anonymous):

You don't want anything under the square root to be negative. So you would set what is under the square root to > or = to 0. And solve.

OpenStudy (anonymous):

pj has it right

OpenStudy (anonymous):

hmm still not understanding

OpenStudy (anonymous):

If you plug zero in for x in \[\sqrt{x^2-1}\] you get \[\sqrt{-1}\] which is imaginary and not good.

OpenStudy (anonymous):

oh ok that wat you meant

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Make sense?